OpenStudy (anonymous):

please help me to solve this question, Express as partial fraction (Repeated linear factors) : 1 / x^3(1-2x)

OpenStudy (luigi0210):

Any ideas?

OpenStudy (luigi0210):

For terms with a power higher than 1, like \(x^2\) for example, you could do this: \(\Large \frac{1}{x^2}=\frac{A}{x}+\frac{B}{x^2}\) ..

OpenStudy (anonymous):

after get equation 1= A(x^2)(1-2x) + B(x)(1-2x) + C(1-2x) + D(x^3), how to find the value of A,B,C,D ? I try to substitute x as 0 , but cant find the ans..

hartnn (hartnn):

I would approach this problem this way : \(\Large \dfrac{1}{x^3 (1-2x)} =\dfrac{-1}{2}\dfrac{1}{x^3 (x-1/2)} \) now for \(\dfrac{1}{x^3 (x-1/2)}\) we can decompose it like \(\Large \dfrac{1}{x^3 (x-1/2)} = \dfrac{A}{x}+\dfrac{B}{x^2}+\dfrac{C}{x^3}+\dfrac{D}{x-1/2}\) now make the common denominator and try to find A,B,C,D values et me know if you get stuck...

OpenStudy (anonymous):

yeah,then i get equation 1 = A1= A(x^2)(1-2x) + B(x)(1-2x) + C(1-2x) + D(x^3), i stuck in finding the value of A,B,C,D.. is it need to expand the equation first? how?

hartnn (hartnn):

ok, so you want to stick with 1-2x anyways, you can find D and C easily plug in x= 0 in the equation and you will get C plug in x= 1/2 in the equation and you will get D try :)

hartnn (hartnn):

for A and B, yes, you need to expand by distributing. then compare the co-efficients....if you don't know how, first find C,D which are easy to get.

OpenStudy (anonymous):

i see.. so the value of C=1, D=8?

hartnn (hartnn):

those are correct :)

OpenStudy (anonymous):

then substitute value of C, D into the equation?

hartnn (hartnn):

that would simplify things, yes. and start distributing in the same step

hartnn (hartnn):

\(1= A(x^2)(1-2x) + B(x)(1-2x) + C(1-2x) + D(x^3) \\ 1=Ax^2 -2Ax^3 +Bx-2Bx+C-2Cx+Dx^3\) \(1=Ax^2 -2Ax^3 +Bx-2Bx+1-2x+8x^3\) \(0=Ax^2 -2Ax^3 +Bx-2Bx-2x+8x^3\)

hartnn (hartnn):

typing error : \(1=Ax^2−2Ax^3+Bx−2Bx^2+1−2x+8x^3\)

hartnn (hartnn):

do you get what i did there ?

OpenStudy (anonymous):

aaa i see..yes i get it, then, to find A just subs x as 0?

hartnn (hartnn):

you already plugged in x=0 to get C, now if you again plug in x=0 you get 0=0 :P so now tell me the co-efficient of \(\Large x^3 \) on right side ??

OpenStudy (anonymous):

ohh i see hehe. erm 2B?

OpenStudy (anonymous):

ehh 2A

hartnn (hartnn):

no...what is before x^3 ?? example in, \(4Ay^2 +2y+By^2\) the co-efficient of y^2 is 4A+B so here, only 2A ?? or someting else too ?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

sry i hv eye problem, -2A , -2B, 8

OpenStudy (anonymous):

not 2B :P

hartnn (hartnn):

its ok not -2B, its with x^2, not x^3 so, co-efficient of x^3 on right side is \(-2A +8\) agree?? and on left its 0 so -2A+8=0 you'll get A from here

hartnn (hartnn):

then compare the co-efficients of x^2 on both sides, and you will get B ask if any more doubts then.

OpenStudy (anonymous):

erm why at the left side is 0, not 1? bit blur ;P

hartnn (hartnn):

\(1=Ax^2 -2Ax^3 +Bx-2Bx^2+1-2x+8x^3\) Subtracting 1 from both sides \(0=Ax^2 -2Ax^3 +Bx-2Bx^2-2x+8x^3\) got this step ?

hartnn (hartnn):

and \(0 = 0+0x+0x^2 +0x^3+....\)

hartnn (hartnn):

so \(0x^3 = -2Ax^3 +8x^3 = (-2A+8)x^3\) so -2A+8= 0 hope this makes everything clear :)

OpenStudy (anonymous):

ohhh now i clearly see.. like it.. ;P

hartnn (hartnn):

similarly compare co-efficient of x^2 from both sides, and u get B too

OpenStudy (anonymous):

so A=1/4, B=1/8?

hartnn (hartnn):

how did u get A =1/4 ? :O

OpenStudy (anonymous):

ohh, 4?

hartnn (hartnn):

yes, now correct :)

OpenStudy (anonymous):

and B=2?

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

thank you very much helping me :)

hartnn (hartnn):

welcome ^_^ happy to help :)

hartnn (hartnn):

you seem to be new here, so \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

yes i am, thank you ^^